3. The position of an object as a function of time in x-and y-coordinate system is given by n-( e k-cy where a 2m.b-5ms, and c- 10 m Find the (a) s- and (b) y components of instantaneous velocity at (e) speed of the objoct at r 3 s (d) total displacement of the object during 3
Answer #1

Given

r(t)=(at^{2}+c)widehat{i}+(bt-c)widehat{j}

where a=2 m/s2 ,b=5 m/s ,c=10

V(t)=dr(t)/dt

V(t) = (20t)i-bj

a)

at t=3 s

Vx=2*2*3 =12 m/s2

b)

Vy=b=5 m/s

c)

V=sqrt[122+52]=13 m/s

d)

displacement

x=Vt =13*3 =39 m